If dy/dx = y − x, y(0) = 1 2 . Use Modified Euler’s method with h = 0.1 to obtain an approximation to y(0.2).


1
Expert's answer
2022-02-24T05:55:42-0500

Solution:

Given "y^{\prime}=y-x, y(0)=0.5, h=0.1, y(0.2)=?"

Here, "x_{0}=0, y_{0}=0.5, h=0.1, x_{n}=0.2"

"\begin{aligned}

&y^{\prime}=y-x \\

&\therefore f(x, y)=y-x

\end{aligned}"

Modified Euler method

"y_{m+1}=y_{m}+h f\left(x_{m}+\frac{1}{2} h, y_{m}+\frac{1}{2} h f\left(x_{m}, y_{m}\right)\right)



\\x_{0}+\frac{1}{2} h=0+\frac{0.1}{2}=0.05



\\y_{0}+\frac{1}{2} h f\left(x_{0}, y_{0}\right)=0.5+\frac{0.1}{2} \cdot 0.5=0.525



\\f\left(x_{0}+\frac{1}{2} h, y_{0}+\frac{1}{2} h f\left(x_{0}, y_{0}\right)=f(0.05,0.525)=0.475\right."

"\begin{aligned}

&y_{1}=y_{0}+h f\left(x_{0}+\frac{1}{2} h, y_{0}+\frac{1}{2} h f\left(x_{0}, y_{0}\right)\right)=0.5+0.1 \cdot 0.475=0.5475 \\

&\therefore y(0.1)=0.5475

\end{aligned}"

Again taking "\left(x_{1}, y_{1}\right)"  in place of "\left(x_{0}, y_{0}\right)"  and repeat the process

"\begin{aligned}

&f\left(x_{1}, y_{1}\right)=f(0.1,0.5475)=0.4475 \\

&x_{1}+\frac{1}{2} h=0.1+\frac{0.1}{2}=0.15 \\

&y_{1}+\frac{1}{2} h f\left(x_{1}, y_{1}\right)=0.5475+\frac{0.1}{2} \cdot 0.4475=0.5699 \\

&f\left(x_{1}+\frac{1}{2} h, y_{1}+\frac{1}{2} h f\left(x_{1}, y_{1}\right)=f(0.15,0.5699)=0.4199\right. \\

&y_{2}=y_{1}+h f\left(x_{1}+\frac{1}{2} h, y_{1}+\frac{1}{2} h f\left(x_{1}, y_{1}\right)\right)=0.5475+0.1 \cdot 0.4199=0.5895 \\

&\therefore y(0.2)=0.5895

\end{aligned}"